456448454/5000 + 93352662/10000
Found 2 solutions by richard1234, MathLover1:
Answer by richard1234(7193)

Other tutor subtracted, not added.

456448454/5000 + 93352662/10000 = (2*456448454 + 93352662)/10000 = 100624957/10000 = 100624.57

And I'm not wrong nor am I an idiot, despite your little "thank you" message.

Answer by MathLover1(20849)